# A Second Main Theorem for Entire Curves Intersecting Three Conics

### v1

@@ 2.1. Logarithmic Jet Bundles and Invariant Differential Forms
Throughout this paper, we only concern the situation $k=1$ or $2$. In the case that $\dim_{\mathbb{C}}X=2$, the bundle of invariant logarithmic jet differentials of order $2$ admits a filtration [Demailly1997]:
@@
where the logarithmic conincal line bundle $\overline{K}_{X}\coloneqq K_{X}\otimes\mathcal{O}_{X}(D)$. This filtration corresponds to the expression of an invariant logarithmic jet differential $P(f^{\prime},f^{\prime\prime})$ outside $D$ as:
@@ 2.2. Logarithmic Demailly-Semple Tower
where $T_{X_{1}/X_{0}}$ is the relative tangent bundle of the fibration $\pi_{1,0}\colon X_{1}\rightarrow X_{0}$. And there is a relative version of the Euler sequence of the tangent bundle of the fiber $\mathbb{P}(V_{0,x})$:
@@
Let $f\colon\mathbb{D}_{R}\rightarrow X_{0}$ be a nonconstant holomorphic map from the disc $\mathbb{D}_{R}=\{z\in\mathbb{C}:|z|<R\}$. To define its liftings to the first level $X_{1}$ of the logarithmic Demailly–Semple tower, we consider the following two cases for $f^{\prime}(t)\in T_{X_{0},f(t)}$:
@@ 2.3. A Bogomolov-type Vanishing Theorem
Later in Sections 3 and 4, to establish the existence of negatively twisted logarithmic $1$-jet, and $2$-jet, differentials via Riemann–Roch, it suffices to verify the vanishing of the cohomology groups $H^{2}\big(\mathbb{P}^{2},S^{m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big)$ and $H^{2}\big(\mathbb{P}^{2},E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big)$, respectively. Using Serre duality and the filtration structure in (2.3), this problem reduces to a Bogomolov-type vanishing theorem for a log pair $(X,D)$ of log general type, where $X$ is a smooth projective surface and $D$ is a reduced simple normal crossing divisor.
@@
4. Twisted Logarithmic $2$-Jet Differentials and Threshold
@@
To apply Siu’s strategy of slanted vector fields, we follow the argument of Siu [Siu2004, pp. 558–559] to extend a negatively twisted logarithmic $2$-jet differential as a holomorphic family, invoking the following standard result in algebraic geometry:
@@
    For each parameter $a\in U^{m,t}$, the corresponding $\mathcal{C}_{a}$ forms a simple normal crossing divisor;
@@
$\omega_{0}\in H^{0}\big(X,E_{2,M}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C}_{a_{0}})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-T)\big).$
Next, we decompose the zero locus $\{\omega_{0}=0\}$ in the second level $X_{2,a_{0}}$ of the Demailly-Semple tower of $(\mathbb{P}^{2},\mathcal{C}_{a_{0}})$ into irreducible components. By the same argument as Demailly and El Goul [Demailly-Elgoul2000, p. 532], [Elgoul2003, Lemma 1.3.2], we can extract an irreducible and reduced component $Z$, which in fact “corresponds” to some irreducible negatively twisted invariant logarithmic $2$-jet differential
(4.3) $\omega_{1}\in H^{0}\big(X,E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C}_{a_{0}})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big),\quad\text{with}\ \tfrac{t}{m}\geqslant\tfrac{T}{M}>\tfrac{1}{5},$
@@
Let $u_{1}=\pi_{2,1}^{*}\mathcal{O}_{X_{1,a_{0}}}(1)$, $u_{2}=\mathcal{O}_{X_{2,a_{0}}}(1)$ and $h=\mathcal{O}_{\mathbb{P}^{2}}(1)$. Then the zero divisor $Z_{\omega_{0}}$ of $\omega_{0}$ on $X_{2,a_{0}}$ is linearly equivalent to
$Mu_{2}-T\pi_{2,0}^{*}h\in\operatorname{Pic}\left(X_{2,a_{0}}\right).$
Let $Z_{\omega_{0}}=\sum p_{j}Z_{j}$ be the irreducible decomposition of the divisor $Z_{\omega_{0}}$, with each $Z_{j}$ irreducible and reduced. Since $\operatorname{Pic}(X_{2,a_{0}})\cong\operatorname{Pic}(\mathbb{P}^{2})\oplus\mathbb{Z}u_{1}\oplus\mathbb{Z}u_{2}$ and $\operatorname{Pic}(\mathbb{P}^{2})\cong\mathbb{Z}$, each component satisfies
@@
whose zero divisor is $Z_{j}+b_{1,j}\Gamma_{2}$. Since both $Z_{j}$ and $\Gamma_{2}$ are irreducible and reduced, $\omega_{j}$ can not be factorized into a product of other logarithmic $2$-jet differentials.
@@ 5. Generating Independent Twisted Jet Differentials via Slanted Vector Fields
To construct an additional family of independent negatively twisted jet differentials, we employ Siu’s slanted vector field method [Siu2004]. This approach generalizes to $k$-jet spaces the original technique developed for $1$-jet spaces by Clemens [Clemens1986], Ein [Ein1988], and Voisin [Voisin1996] in their study of rational curves.
@@
This section adapts the standard technique of Siu’s slanted vector fields—twisted by $\mathcal{O}_{\mathbb{P}^{2}}(7)$—to the case of three generic conics in $\mathbb{P}^{2}$. The content is entirely classical and no originality is claimed.
@@ 5.1. Logarithmic Slanted Vector Fields for Three Conics in $\mathbb{P}^{2}$
where for each $1\leqslant\lambda\leqslant 3$, the hypersurface $\mathcal{Y}_{\lambda}$ is the universal family of degree $d_{\lambda}$ hypersurfaces in $\mathbb{P}^{5}$, given explicitly by $\mathcal{Y}_{\lambda}:=\left\{A_{\mathbf{0}}^{\lambda}Z_{2+\lambda}^{d_{\lambda}}+\sum_{|\alpha|=d_{\lambda}}A_{\alpha}^{\lambda}Z^{\alpha}=0\right\}\subset\mathbb{P}^{5}\times\mathbb{P}^{N_{d_{\lambda}}+1}$. Here, the space $\mathbb{P}^{5}$ has homogeneous coordinates $Z:Z_{3}:Z_{4}:Z_{5}]=[Z_{0}:\cdots:Z_{5}]$, and the coefficient space $\mathbb{P}^{N_{d_{\lambda}}+1}$ is equipped with homogeneous coordinates $A^{\lambda}:A_{\mathbf{0}}^{\lambda}]=[\ldots:A_{\alpha}^{\lambda}:\ldots:A_{\mathbf{0}}^{\lambda}]_{|\alpha|=d_{\lambda}}$.
@@
Now we consider $\mathcal{Y}^{*}\coloneqq\mathcal{Y}\setminus\left(\cup_{\lambda=1}^{3}\left(\{A_{\mathbf{0}}^{\lambda}=0\}\cup\{A^{\lambda}=\mathbf{0}\}\right)\right)$ on which the restricted forgetful map $\pi|_{\mathcal{Y}^{*}}\colon\mathcal{Y}^{*}\rightarrow\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$ is well-defined according to the containment relation (5.2). Furthermore, we obtain the identification of the preimage $\left(\pi|_{\mathcal{Y}^{*}}\right)^{-1}(\mathcal{X})=\{Z_{3}Z_{4}Z_{5}=0\}=:\mathcal{D}$. This shows that $\pi$ induces a logarithmic morphism from $\left(\mathcal{Y}^{*},\mathcal{D}\right)$ to $\big(\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}},\mathcal{X}\big)$, which consequently gives rise to a dominant map between jet spaces:
@@
and $\mathrm{pr}_{1}$ (resp. $\mathrm{pr}_{2}$) is the projection from $\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$ to $\mathbb{P}^{2}$ (resp. ${\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$).
@@ 5.2. Producing an Independent $2$-Jet via Siu’s Slanted Vector Fields of Pole Order 77
$P\colon\overline{J_{2}^{v}}\left(\mathbb{P}^{2}\times{\textstyle\prod}_{{\lambda}=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}\right)\Big|_{\mathrm{pr}_{2}^{-1}\left(U\right)}\longrightarrow p^{*}\mathrm{pr}_{1}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-t)|_{\mathrm{pr}_{2}^{-1}\left(U\right)},$
@@ 6.1. More Riemann-Roch Computations
In the preceding Section 4, we show the existence of a nonzero negatively twisted invariant logarithmic $2$-jet differential
@@
$\mathcal{O}_{X_{2}}(b_{1},b_{2})\coloneqq\pi_{2,1}^{*}\mathcal{O}_{X_{1}}(b_{1})\otimes\mathcal{O}_{X_{2}}(b_{2}).$
By the same argument as Demailly–El Goul [Demailly-Elgoul2000, Proposition 3.4], one can show that $\mathcal{O}_{X_{2}}(2,1)|_{Z}$ is big when $t<m$. Now, our goal is to determine an effective small rational number $\tau>0$ such that the $\mathbb{Q}$-line bundle $\big(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau)\big)|_{Z}$ remains big. This will be achieved through Riemann-Roch computations and the following proposition observed by Demailly [Demailly1997, p. 73].
@@
Let $x\in X_{0}$ be an arbitrary point and denote by $F_{1}\cong\mathbb{P}^{1}$ the fiber of $\pi_{1,0}:X_{1}\to X_{0}$ over $x$. From the construction of the logarithmic Demailly–Semple tower (Subsection 2.2) we know that the fiber $F_{2}$ of $\pi_{2,0}$ over $x$ satisfies
@@
If $0<\frac{t}{m}<1$ and $0\leqslant\tau<\tau_{1}(\frac{t}{m})$, where
@@
Thus when $0<\frac{t}{m}<1$ and $0\leqslant\tau<\tau_{1}(\frac{t}{m})$, we have $(2u_{1}+u_{2}-\tau\pi_{2,0}^{*}h)^{3}\cdot Z>0$. Applying the Riemann-Roch Theorem then gives, for sufficiently divisible integer $p\gg 1$, that:
@@
$\mathcal{O}_{X_{2}}(-Z)\otimes\left(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau)\right)^{p}\cong\mathcal{O}_{X_{2}}(2p-b_{1},\,p-b_{2})\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-(\tau p+t)),$
@@
Thus, by Leray’s spectral sequence, we conclude that
@@
$\cong H^{3}\big(\mathbb{P}^{2},(\pi_{2,0})_{*}(\mathcal{O}_{X_{2}}(-Z)\otimes(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p})\big)=0,$
for all sufficiently divisible $p\gg 1$. This establishes (6.6), which completes the proof. ∎

### v2

@@ 2.1. Logarithmic Jet Bundles and Invariant Differential Forms
Throughout this article, we only concern the situation $k=1$ or $2$. In the case that $\dim_{\mathbb{C}}X=2$, the bundle of invariant logarithmic jet differentials of order $2$ admits a filtration structure with the grading [Demailly1997]:
@@
where $\overline{K}_{X}$ is the logarithmic canonical line bundle $K_{X}\otimes\mathcal{O}_{X}(D)$. This filtration corresponds to the expression of an invariant logarithmic $2$-jet differential $P(f^{\prime},f^{\prime\prime})$ outside $D$ as:
@@ 2.2. Logarithmic Demailly-Semple Tower
where $T_{X_{1}/X_{0}}$ is the relative tangent bundle of the fibration $\pi_{1,0}\colon X_{1}\rightarrow X_{0}$. Moreover, there is a relative version of the Euler sequence of the tangent bundle of the fiber $\mathbb{P}(V_{0,x})$:
@@
Let $f\colon\mathbb{D}_{R}\rightarrow X_{0}$ be a nonconstant holomorphic map from the disc $\mathbb{D}_{R}=\{z\in\mathbb{C}:|z|<R\}$ of radius $R\in\mathbb{R}_{+}\cup\{\infty\}$. To define its liftings to the first level $X_{1}$ of the logarithmic Demailly–Semple tower, we consider the following two cases for $f^{\prime}(t)\in T_{X_{0},f(t)}$:
@@ 2.3. A Bogomolov-type Vanishing Theorem
In the subsequent Sections 3 and 4, the existence of negatively twisted logarithmic $1$-jet and $2$-jet differentials will be established via Riemann–Roch calculations. This reduces to verifying the vanishing of the corresponding second cohomology groups:
$H^{2}\big(\mathbb{P}^{2},S^{m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big)\quad\text{and}\quad H^{2}\big(\mathbb{P}^{2},E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big),$
respectively. A key tool in proving these vanishing results is the following Bogomolov-type vanishing theorem.
@@
4. Twisted Invariant Logarithmic $2$-Jet Differentials and Threshold
@@
To apply Siu’s strategy of slanted vector fields, we follow the argument of Siu [Siu2004, pp. 558–559] to extend a negatively twisted invariant logarithmic $2$-jet differential as a holomorphic family, invoking the following standard result in algebraic geometry:
@@
    For each parameter $a\in U^{m,t}$, the corresponding $\mathcal{C}_{a}$ forms a simple normal crossing divisor.
@@
$\omega_{0}\in H^{0}\big(\mathbb{P}^{2},E_{2,M}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C}_{a_{0}})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-T)\big).$
Next, we decompose the zero locus $\{\omega_{0}=0\}$ in the second level $X_{2,a_{0}}$ of the logarithmic Demailly-Semple tower of $(\mathbb{P}^{2},\mathcal{C}_{a_{0}})$ into irreducible components. By the same argument as Demailly and El Goul [Demailly-Elgoul2000, p. 532], [Elgoul2003, Lemma 1.3.2], we can extract an irreducible and reduced component $Z$, which in fact “corresponds” to some irreducible negatively twisted invariant logarithmic $2$-jet differential
(4.3) $\omega_{1}\in H^{0}\big(\mathbb{P}^{2},E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C}_{a_{0}})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big),\quad\text{with}\ \tfrac{t}{m}\geqslant\tfrac{T}{M}>\tfrac{1}{5},$
@@
Let $u_{1}=\pi_{2,1}^{*}\mathcal{O}_{X_{1,a_{0}}}(1)$, $u_{2}=\mathcal{O}_{X_{2,a_{0}}}(1)$ and $h=\mathcal{O}_{\mathbb{P}^{2}}(1)$. Then the zero divisor $Z_{\omega_{0}}=\{\omega_{0}=0\}\subset X_{2,a_{0}}$ is linearly equivalent to $Mu_{2}-T\pi_{2,0}^{*}h$ in the Picard group
$\operatorname{Pic}(X_{2,a_{0}})\cong\operatorname{Pic}(\mathbb{P}^{2})\oplus\mathbb{Z}u_{1}\oplus\mathbb{Z}u_{2}\cong\mathbb{Z}^{3}.$
Let $Z_{\omega_{0}}=\sum p_{j}Z_{j}$ be the irreducible decomposition of the divisor $Z_{\omega_{0}}$, with each $Z_{j}$ irreducible and reduced. Decompose each component $Z_{j}$ linear equivalently in the Picard group $\operatorname{Pic}(X_{2,a_{0}})$ as
@@
whose zero divisor is $Z_{j}+b_{1,j}\Gamma_{2,a_{0}}$. Since both $Z_{j}$ and $\Gamma_{2,a_{0}}$ are irreducible and reduced, $\omega_{j}$ can not be factorized into a product of other logarithmic $2$-jet differentials.
@@ 5. Generating Independent Twisted Jet Differentials via Slanted Vector Fields
To construct an additional family of independent negatively twisted jet differentials, we employ Siu’s method of slanted vector fields [Siu2004]. This approach generalizes to $k$-jet spaces the original technique developed for $1$-jet spaces by Clemens [Clemens1986], Ein [Ein1988], and Voisin [Voisin1996] in their study of rational curves.
@@
This section adapts the standard technique of Siu’s slanted vector fields to the case of three generic conics in $\mathbb{P}^{2}$. The content is entirely classical and no originality is claimed.
@@ 5.1. Logarithmic Slanted Vector Fields for Three Conics in $\mathbb{P}^{2}$
where for each $1\leqslant\lambda\leqslant 3$, the hypersurface $\mathcal{Y}_{\lambda}$ is the universal family of degree $d_{\lambda}$ hypersurfaces in $\mathbb{P}^{5}$, given explicitly by $\mathcal{Y}_{\lambda}:=\left\{A_{\mathbf{0}}^{\lambda}Z_{2+\lambda}^{d_{\lambda}}+\sum_{|\alpha|=d_{\lambda}}A_{\alpha}^{\lambda}Z^{\alpha}=0\right\}\subset\mathbb{P}^{5}\times\mathbb{P}^{N_{d_{\lambda}}+1}$. Here, the space $\mathbb{P}^{5}$ has homogeneous coordinates $Z:Z_{3}:Z_{4}:Z_{5}]=[Z_{0}:\cdots:Z_{5}]$, and the parameter space $\mathbb{P}^{N_{d_{\lambda}}+1}$ is equipped with homogeneous coordinates $A^{\lambda}:A_{\mathbf{0}}^{\lambda}]=[\ldots:A_{\alpha}^{\lambda}:\ldots:A_{\mathbf{0}}^{\lambda}]_{|\alpha|=d_{\lambda}}$.
@@
Now we consider $\mathcal{Y}^{*}\coloneqq\mathcal{Y}\setminus\left(\cup_{\lambda=1}^{3}\left(\{A_{\mathbf{0}}^{\lambda}=0\}\cup\{A^{\lambda}=\mathbf{0}\}\right)\right)$ on which the restricted forgetful map
$\pi|_{\mathcal{Y}^{*}}\colon\mathcal{Y}^{*}\rightarrow\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$
is well-defined according to the containment relation (5.2). Furthermore, we obtain the identification of the preimage $\left(\pi|_{\mathcal{Y}^{*}}\right)^{-1}(\mathcal{X})=\{Z_{3}Z_{4}Z_{5}=0\}=:\mathcal{D}$. This shows that $\pi$ induces a logarithmic morphism from $\left(\mathcal{Y}^{*},\mathcal{D}\right)$ to $\big(\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}},\mathcal{X}\big)$, which consequently gives rise to a dominant map between jet spaces:
@@
and $\mathrm{pr}_{1}$ (resp. $\mathrm{pr}_{2}$) is the projection from $\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$ to $\mathbb{P}^{2}$ (resp. ${\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$). ∎
@@ 5.2. Producing an Independent $2$-Jet via Siu’s Slanted Vector Fields of Pole Order 77
$P\colon\overline{J_{2}^{v}}\left(\mathbb{P}^{2}\times{\textstyle\prod}_{{\lambda}=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}\right)\Big|_{\mathrm{pr}_{2}^{-1}\left(U\right)}\longrightarrow p^{*}\mathrm{pr}_{1}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-t)|_{\mathrm{pr}_{2}^{-1}\left(U\right)}.$
@@ 6.1. More Riemann-Roch Computations
In Proposition 4.4, we show the existence of a nonzero negatively twisted invariant logarithmic $2$-jet differential
@@
$\mathcal{O}_{X_{2}}(b_{1},b_{2})\coloneqq\pi_{2,1}^{*}\mathcal{O}_{X_{1}}(b_{1})\otimes\mathcal{O}_{X_{2}}(b_{2}),\quad b_{1},b_{2}\in\mathbb{Z}.$
Using the same argument as in Demailly–El Goul [Demailly-Elgoul2000, Proposition 3.4] — an idea which originates in the work of Miyaoka [Miyaoka] and Lu–Yau [Lu-Yau] — one can show that $\mathcal{O}_{X_{2}}(2,1)|_{Z}$ is big, provided that $t<m$. Now, our goal is to determine an effective small rational number $\tau>0$ such that the $\mathbb{Q}$-line bundle $\big(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau)\big)|_{Z}$ remains big. This will be achieved through Riemann-Roch computations and the following proposition observed by Demailly [Demailly1997, p. 73].
@@
Let $x\in X_{0}$ be an arbitrary point and denote by $F_{1}\cong\mathbb{P}^{1}$ the fiber of $\pi_{1,0}:X_{1}\to X_{0}$ over $x$. From the construction of the logarithmic Demailly-Semple tower (Subsection 2.2) we know that the fiber $F_{2}$ of $\pi_{2,0}$ over $x$ satisfies
@@
If $0<\frac{t}{m}<1$ and $0\leqslant\tau<\tau_{1}(\frac{t}{m}),\tau\in\mathbb{Q}$, where
@@
Thus when $0<\frac{t}{m}<1$ and $0\leqslant\tau<\tau_{1}(\frac{t}{m})$, we have $(2u_{1}+u_{2}-\tau\pi_{2,0}^{*}h)^{3}\cdot Z>0$. Riemann-Roch Theorem then gives, for sufficiently divisible integer $p\gg 1$, that:
@@
$\mathcal{O}_{X_{2}}(-Z)\otimes\left(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau)\right)^{p}\cong\mathcal{O}_{X_{2}}(2p-b_{1},\,p-b_{2})\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(t-\tau p),$
@@
Thus, by Leray’s spectral sequence, we conclude that for all sufficiently divisible $p\gg 1$:
@@
[…]
$=\$$0\qquad\text{[because\,\,\, $\dim\mathbb{P}^{2}=2<3$]}.$
This establishes (6.6), which completes the proof. ∎